Please help me solve this question.
Show that the real solution of the equation ax2+bx+c=0 are the negative of the real solution of the equation ax2-bx+c=0. Assume that b2-4ac ≥ 0.
Thank you

Answer by stanbon(75887)

Show that the real solution of the equation ax2+bx+c=0 are the negative of the real solution of the equation ax2-bx+c=0. Assume that b2-4ac ≥ 0.
----
Use the Quadratic Formula:
ax^2 + bx + c = 0
has solutions: x = [-b+-sqrt(b^2-4ac)]/(2a)
-------
ax^2-bx + c = 0
has solutions: x = [b +- sqrt(b^2-4ac)]/(2a)
-------
[-b+-sqrt(b^2-4ac)]/2a ? [b +- sqrt(b^2-4ac)]/(2a)
-----
-b/(2a) ? b/(2a)
----
-b/(2a) is the negative of b/(2a)
====================================
Cheers,
Stan H.